OpenStudy (marsxtc):

Implicit Differentiation (x+y)^3 = x^3+y^3 find dy/dx at points (-1,1)

OpenStudy (marsxtc):

Im pretty confused. I'm starting out with 3(x+y)^2 * (1+dy/dx) on the left side. Currently redoing it

OpenStudy (marsxtc):

Yeah. I don't know where I'm going with this.

OpenStudy (anonymous):

x=0 y=0

OpenStudy (marsxtc):

Thats not what the questions looking for

OpenStudy (marsxtc):

Thanks for trying :X

OpenStudy (amistre64):

your initial work is fine

OpenStudy (marsxtc):

Currently on the left side I end up with 3(x+y)^2 * ( 1+ dy/dx) = 3x^2(dx/dx) + 3y^2(dy/dx) right side <-- am I going on the right path?

OpenStudy (anonymous):

sorry I just looked at it and didn't think lol :D I guess I need to pay more attention

OpenStudy (amistre64):

consider this, since you are trying this with the d/dx notations: d/dx [f] = df/dx d/dx [y] = dy/dx d/dx [x] = dx/dx, which is why its 1

OpenStudy (amistre64):

youre doing great so far

OpenStudy (marsxtc):

I'm trying to do this step by step, so on the left side after simplifying it I ended up with (3+3dy/dx) (x+y)^2.

OpenStudy (marsxtc):

The goal is to find dy/dx too by the way, at the point of (-1,1)

OpenStudy (amistre64):

if we just let a ' define the derivative, the writeup goes \[[(x+y)^3=x^3+y^3]'\] \[[(x+y)^3]'=[x^3]'+[y^3]'\] \[3(x+y)^2~[(x+y)]'=3x^2x'+3y^2y'\] \[3(x+y)^2~(x'+y')=3x^2x'+3y^2y'\] since x' = dx/dx = 1 \[3(x+y)^2~(1+y')=3x^2+3y^2y'\] the rest is algebra to solve for y'

OpenStudy (amistre64):

you can plug in the xy values at this stage if all the xy parts are confusing :) then algebra the y' out of it ... your call

OpenStudy (marsxtc):

So I plugged in (-1, and 1) for the last equation you gave me ( I got up to that too ) and I got derivative of y = -1. Is that correct? If It is Can you help me with a problem or two more?

OpenStudy (amistre64):

notice that both sides have 3 as a common factor, so instead of trying to distribute it we can divide it off \[\frac{\cancel3(x+y)^2~(1+y')=\cancel3x^2+\cancel3y^2y'}{\cancel3}\] \[(x+y)^2~(1+y')=x^2+y^2y'\] \[(-1+1)^2~(1+y')=1^2+(-1)^2y'\] \[0=1+y'~:~y'=-1\]

OpenStudy (marsxtc):

Yep -1 :) Thank you! Can you help me out with one or two more? I had 10 problems and only these few confused/ made me double guess myself

OpenStudy (amistre64):

i can spot you on 1 more ...

OpenStudy (marsxtc):

So for sqrtxy = x-2y at (4,1)

OpenStudy (marsxtc):

I got, 1/2(xy)^-1/2 * [x dy/dx + y dx/dx] ( for the left side )

OpenStudy (amistre64):

good so far, an eyesore, but good

OpenStudy (marsxtc):

So then I simplified the first term, which I got 1/2sqrtxy then I multiplied it by the derivatives = (x dy/dx + y) / 2sqrt xy

OpenStudy (amistre64):

sqrt(xy) is more acceptable ascii notation; but go on

OpenStudy (marsxtc):

I then started working on the left side where I got 1-2 d(y)/d(x), Im unsure what to do now Should I multiply both sides by 2sqrt(xy)?

OpenStudy (amistre64):

xy' + y ------- = 1 - 2y' 2sqrt(xy) if you want to do more work to it, thats up to you; but if all you needs a a value of y' at (4,1), id just plug and play it 4y' + 1 ------- = 1 - 2y' 2sqrt(4) 4y' + 1 ------- = 1 - 2y' 4 4y' + 1 = 4 - 8y' now its more like a simple algebra problem .....

OpenStudy (marsxtc):

Alright, thanks, but it could eventually start working out if I did cancel out the denominator on the left side right?

OpenStudy (amistre64):

of course, but i find that there comes a point when doing alot of "simplifying" to develop a formula for y' is just pointless is all

OpenStudy (marsxtc):

Haha thanks so much! Yeah I've been taught to just simplify to the end! Thank you for all your help!

OpenStudy (amistre64):

youre welcome :) putting all that energy and time into simplifying when a text might only need a final number .... potato potato i spose lol

OpenStudy (amistre64):

*a test....

OpenStudy (marsxtc):

Just a question, are you a real life math teacher or?